OpenStudy (anonymous):

Find the general form of the equation of the the circle. Center at the point (-4, -3); containing the point (-3, 3)

OpenStudy (anonymous):

general form is \[(x-h)^2+(y-k)^2=r^2\] so you will have \[(x+4)^2+(y+3)^2=r^2\] and now you just need \[r^2\]

OpenStudy (anonymous):

to find it find the square of the distance between these point. i t is \[r^2=(-4+3)^2+(-3-3)^2=1+36=37\] so y our answer is \[(x+4)^2+(y+3)^2=37\]